OpenStudy (narissa):

what is the value of the expression 1/5^-5

OpenStudy (whovianchick):

3125, since a negative exponent is in the denominator, you have to make it positive and move it to the numerator, or 1*5^5/1= 5^5 = 3125

OpenStudy (narissa):

im not sure if that's right

OpenStudy (whovianchick):

Why not?

OpenStudy (whovianchick):

\[1\div5^5\]

OpenStudy (whovianchick):

wait I typed that last one wrong, imagine the little five is negative

OpenStudy (whovianchick):

You move it to the numerator, and make it positive so it's \[1\times5^5\]

OpenStudy (whovianchick):

5^5= 3125